OpenStudy (anonymous):

Differentiate w.r.t. x, \large y=\ln \frac{ \sqrt[3]{2x^3-14}}{(1-x^2)^5}

OpenStudy (anonymous):

\large y=\ln \frac{ \sqrt[3]{2x^3-14}}{(1-x^2)^5}\[y=In \sqrt[3]{2x^3-14}/(1-x^2)^5 \]

OpenStudy (anonymous):

\[y ' =x^2/x^3-7 - 10x/1-x^2 \]

OpenStudy (experimentx):

??

OpenStudy (anonymous):

My question is, should we differentiate it more?

OpenStudy (experimentx):

could you revise that question??

OpenStudy (mimi_x3):

\[\large y = \frac{\ln\sqrt[3]{2x^{3}-14}}{(1-x^{2})^{5}} \] it didnt seem like youd differentiated it yet from what you post after it

OpenStudy (anonymous):

yes, I differentiated it. But I have to ask should I differentiate it more? or is that the answer?

OpenStudy (experimentx):

if you need to verify it ...you can wolfram it

OpenStudy (anonymous):

how can I write in wolfram this question? Can you please tell me

OpenStudy (experimentx):

use ()

OpenStudy (experimentx):

(ln(2x^^3 - 14)^(1/3))/(1-x^3)^5

OpenStudy (anonymous):

Thank you both of you for your time and help. :)